# Classification Diffusion Models: Revitalizing Density Ratio Estimation

### v2

@@ Abstract
A prominent family of methods for learning data distributions relies on density ratio estimation (DRE), where a model is trained to classify between data samples and samples from some reference distribution. DRE-based models can directly output the likelihood for any given input, a highly desired property that is lacking in most generative techniques. Nevertheless, to date, DRE methods have struggled to accurately capture the distributions of complex high-dimensional data like images, which led to reduced research attention over the years. In this work we present classification diffusion models (CDMs), a DRE-based generative method that adopts the formalism of denoising diffusion models (DDMs) while making use of a classifier that predicts the level of noise added to a clean signal. Our method is based on an analytical connection that we derive between an MSE-optimal denoiser for white Gaussian noise and a cross-entropy-optimal classifier for predicting the noise level. To the best of our knowledge, our method is the first DRE-based technique that can successfully generate images. Furthermore, it can output the likelihood of any input in a single forward pass, achieving state-of-the-art negative log likelihood (NLL) among methods with this property. Code is available on the project’s webpage.
@@ 1 Introduction
A classical family of methods for learning data distributions relies on the concept of density-ratio estimation (DRE) [45]. DRE techniques transform the unsupervised task of learning the distribution of data into the supervised task of learning to classify between data samples and samples from some reference distribution [15, 4, 34, 7]. These methods have attracted significant research efforts over the years [27, 14, 34, 46], particularly for their inherent capability to directly output the likelihood for any given input. However, to date, they have not succeeded to successfully capture the distribution of complex high-dimensional data, like natural images, and their generative performance was demonstrated only on toy examples or on the simple MNIST handwritten digits dataset [28, 34, 7]. As illustrated in Fig. 1, the state-of-the-art DRE method, telescoping density-ratio estimation (TRE) [34], succeeds in capturing the distribution of MNIST [28], but fails on CIFAR-10 [26].
@@
Our experiments shed light on the reasons why DRE methods have failed to effectively learn complex high-dimensional data to date, and why CDM inherently avoids these challenges. Furthermore, we show that CDM can serve as a more accurate denoiser, in terms of MSE, than a DDM with a similar architecture. This often leads to better FID scores. Representative generated samples are shown in Fig. 1. We hope that our approach will spark new interest in DRE methods and ultimately unlock their full potential.
Figure 1: Samples from CDM (left) of CelebA $\mathbf{64\times 64}$ and CIFAR-10, compared to samples from TRE [34] (right) of MNIST and CIFAR-10. To date, DRE methods have failed to capture the distribution of complex, high-dimensional data like images and have been demonstrated only on toy examples or on the simple MNIST dataset. As can be seen in the rightmost column, TRE, the state-of-the-art DRE method, has failed to capture the distribution of CIFAR-10. CDM is the first DRE-based method that can successfully learn the distribution of images.
@@ 2.1 Density Ratio Estimation
Unfortunately, in practice, this method fails when $p_{d}({\bm{x}})$ and $p_{n}({\bm{x}})$ differ significantly from one another, as is the case when $p_{d}({\bm{x}})$ is the distribution of images and $p_{n}({\bm{x}})$ corresponds to white Gaussian noise. This is because when training a classifier to discriminate between images and noise it can achieve high accuracy without learning meaningful information about images. When this point is reached, the weights of the classifier practically stop updating. Rhodes et al. [34] referred to this issue as the density-chasm problem, and suggested to overcome it by making the classification problem more difficult. To do so, their TRE method uses a sequence of distributions $p_{{\mathbf{x}}_{0}}({\bm{x}}),p_{{\mathbf{x}}_{1}}({\bm{x}}),\ldots,p_{{\mathbf{x}}_{m}}({\bm{x}})$, which are closer to one another, such that $p_{{\mathbf{x}}_{m}}({\bm{x}})$ is the reference distribution and $p_{{\mathbf{x}}_{0}}({\bm{x}})$ is the target distribution. The intermediate distributions $\{p_{{\mathbf{x}}_{i}}({\bm{x}})\}_{1}^{m-1}$ do not have to be known; the only requirement is that it would be possible to sample from them. For example, ${\mathbf{x}}_{i}$ can be defined as $\smash{{\mathbf{x}}_{i}=\sqrt{\bar{\alpha}_{i}}{\mathbf{x}}_{0}+\sqrt{1-\bar{\alpha}_{i}}{\mathbf{x}}_{m}}$, where ${\mathbf{x}}_{0}\sim p_{{\mathbf{x}}_{0}}$, ${\mathbf{x}}_{m}\sim p_{{\mathbf{x}}_{m}}$, and $\bar{\alpha}_{i}$ is a sequence that decreases from 1 to 0. Then, using (1), each ratio $p_{{\mathbf{x}}_{i}}({\bm{x}})/p_{{\mathbf{x}}_{i+1}}({\bm{x}})$ can be extracted by training a binary classifier to distinguish between samples from $p_{{\mathbf{x}}_{i}}({\bm{x}})$ and $p_{{\mathbf{x}}_{i+1}}({\bm{x}})$, and the ratio between the target and the reference distributions can be calculated using
@@
While this method overcomes the density-chasm-problem for each pair of consecutive distributions, it still fails in learning the distribution of datasets that are more complicated than MNIST, as illustrated in Fig. 1. This is because each ratio $p_{{\mathbf{x}}_{i}}({\bm{x}})/p_{{\mathbf{x}}_{i+1}}({\bm{x}})$ is extracted from a binary classifier trained only on inputs ${\bm{x}}$ from the distributions $p_{{\mathbf{x}}_{i}}$ and $p_{{\mathbf{x}}_{i+1}}$. For instance, the classifier producing the ratio $p_{{\mathbf{x}}_{0}}({\bm{x}})/p_{{\mathbf{x}}_{1}}({\bm{x}})$ is trained on inputs close to the real data, while the one producing $p_{{\mathbf{x}}_{m-1}}({\bm{x}})/p_{{\mathbf{x}}_{m}}({\bm{x}})$ is trained on inputs close to the reference distribution. This can lead to a mismatch between training and test time, since at inference, all the ratios are evaluated at the same input ${\bm{x}}$. Our method is also based on a classification objective, however it avoids this problem by employing an additional loss, which is based on our main result (Theorem 3.1).
@@ 2.2 Denoising Diffusion Models
DDMs, first introduced by Sohl-Dickstein et al. [37], Ho et al. [19], are a class of generative models that sample from a learned target distribution by gradually denoising white Gaussian noise. More formally, DDMs generate samples by attempting to reverse a forward diffusion process with $T$ steps that starts from a data point ${\mathbf{x}}_{0}$ and evolves as ${\mathbf{x}}_{t}=\sqrt{\alpha_{t}}{\mathbf{x}}_{t-1}+\sqrt{1-\alpha_{t}}\tilde{\varepsilon}_{t}$, $t=1,\ldots,T$, where $\{\tilde{\varepsilon}_{t}\}$ are iid standard Gaussian vectors. Samples along this forward diffusion process can be equivalently expressed as
@@
and covariance $\sigma_{t}\mathbf{I}$, where $\varepsilon_{\theta}(\cdot,\cdot)$ is a neural network and $\sigma_{t}$ is a fixed hyperparameter. Training is done by minimizing the ELBO loss, which reduces to a series of MSE terms,
$\mathcal{L}=\sum_{t=1}^{T}\mathbb{E}_{{\mathbf{x}}_{0},{\mathbf{\varepsilon}}_{t}}\left[\lVert\varepsilon_{\theta}({\mathbf{x}}_{t},t)-{\mathbf{\varepsilon}}_{t}\rVert_{2}^{2}\right].$ (5)
@@ 3 Method
Given a sample ${\bm{x}}$ drawn from (7), we are interested in a classifier that outputs the probability vector $(p_{{\textnormal{t}}|\tilde{{\mathbf{x}}}}(0|{\bm{x}}),p_{{\textnormal{t}}|\tilde{{\mathbf{x}}}}(1|{\bm{x}}),\ldots,p_{{\textnormal{t}}|\tilde{{\mathbf{x}}}}(T+1|{\bm{x}}))$, where $p_{{\textnormal{t}}|\tilde{{\mathbf{x}}}}(t|{\bm{x}})=\mathbb{P}({\textnormal{t}}=t|\tilde{{\mathbf{x}}}={\bm{x}})$. As we now show, the denoiser in (6) corresponds to the gradient of this classifier.
@@
Theorem 3.1 suggests that we may train a classifier and use its gradient as a denoiser according to relation (8), as illustrated in Fig. 2, and then apply any desired sampling method (e.g. DDPM, DDIM, etc.). However, as we show in Sec. 4.1, naively training such a classifier with the standard cross-entropy (CE) loss leads to poor results. This is because a classifier may reach low CE loss even without learning the correct probability $p_{{\textnormal{t}}|\tilde{{\mathbf{x}}}}(t|{\bm{x}})$ for any $t$. This phenomenon can be observed in Fig. 3, which illustrates the reason that to date DRE methods failed to capture the distribution of high-dimensional complex data like images. We discuss this in more detail in Sec. 4.1.
@@
To obtain the correct probability $p_{{\textnormal{t}}|\tilde{{\mathbf{x}}}}(t|{\bm{x}})$ for any $t$, we suggest training the classifier with a combination of a CE loss on its outputs, and an MSE loss on its gradient, according to relation (8). The network’s gradient can be efficiently computed using automatic-differentiation. Following Ho et al. [19], we use the same weight for all timesteps in the MSE loss. Our full training algorithm is described in Algorithm 1, where $f_{\theta}({\bm{x}})[t]$ is the model’s $t$-th logit, which serves as an approximation for $\log p_{{\textnormal{t}}|\tilde{{\mathbf{x}}}}(t|{\bm{x}})$ (up to an additive constant that cancels out in the SoftMax operation), and $\smash{F_{\theta}({\bm{x}},t)=f_{\theta}({\bm{x}})[T+1]-f_{\theta}({\bm{x}})[t]}$. The added timesteps, corresponding to entries 00 and $T+1$ of the classifier, are trained only using the CE loss. This is because the prediction of the noise is trivial when ${\textnormal{t}}=T+1$ and meaningless when ${\textnormal{t}}=0$ (since there is no noise). Importantly, this behavior is automatically achieved without any modification of the algorithm. Specifically, in line 4 of the algorithm, $F_{\theta}({\bm{x}}_{T+1},T+1)=0$, and in line 5, $\varepsilon_{\theta}({\bm{x}}_{0},0)=0$, preventing the MSE loss from updating the weights for these timesteps.
Algorithm 2 shows how to generate samples with CDM using the DDPM sampler, while a similar approach can be used with other samplers. Note that each step $t$ in DDPM sampling using CDM is given by
@@ 3.1 Exact Likelihood Calculation in a Single Step
$p_{{\mathbf{x}}_{t}}({\bm{x}})=\frac{p_{\textnormal{t}}(T+1)}{p_{\textnormal{t}}(t)}\,\frac{p_{{\textnormal{t}}|\tilde{{\mathbf{x}}}}(t|{\bm{x}})}{p_{{\textnormal{t}}|\tilde{{\mathbf{x}}}}(T+1|{\bm{x}})}\,\mathcal{N}({\bm{x}};0,\mathbf{I}).$ (10)
Note that the first term in (10) only depends on the pre-selected probability mass function $p_{\textnormal{t}}$ (which we choose to be uniform in our experiments), and the second term can be obtained from the $t$-th and $(T+1)$-th entries of the vector at the output of the classifier (after applying SoftMax). This implies that we can calculate the likelihood of any given image ${\bm{x}}$ w.r.t. to the density $p_{{\mathbf{x}}_{t}}$ of noisy images for any noise level $t$. In particular, $p_{{\mathbf{x}}_{0}}$ is the density of clean images. Thus, by choosing $t=0$ we can calculate the log likelihood of clean images. In App. D we present a simple validation of our efficient likelihood computation by applying it to toy examples with known densities, allowing us to compute the analytical likelihood and verify that our computations align with theoretical expectations.
Figure 3: Comparison of the log probability of a single prediction between a model trained only using CE and a model trained using CE and MSE. Since the SoftMax operator is invariant to an additive factor, we subtract the maximal value from the vector (i.e., $f_{\theta}({\bm{x}}_{t})\leftarrow f_{\theta}({\bm{x}}_{t})-\max(f_{\theta}({\bm{x}}_{t}))$) for visualization. We utilize the connection we developed between the optimal classifier and the MMSE denoiser to incorporate the MSE loss in DRE training, as depicted in Algorithm 1. As evident, without considering MSE, the prediction accuracy of the classifier is limited to the vicinity of the correct label, unlike the model trained using both CE and MSE, which yields accurate predictions globally. The essence of an optimal classifier lies in its capability to predict the correct probability vector for all entries, rather than solely for the correct label. As can be seen, this necessitates the incorporation of the MSE loss.

### v3

@@ Abstract
A prominent family of methods for learning data distributions relies on density ratio estimation (DRE), where a model is trained to classify between data samples and samples from some reference distribution. DRE-based models can directly output the likelihood for any given input, a highly desired property that is lacking in most generative techniques. Nevertheless, to date, DRE methods have failed in accurately capturing the distributions of complex high-dimensional data, like images, and have thus been drawing reduced research attention in recent years. In this work we present classification diffusion models (CDMs), a DRE-based generative method that adopts the formalism of denoising diffusion models (DDMs) while making use of a classifier that predicts the level of noise added to a clean signal. Our method is based on an analytical connection that we derive between the MSE-optimal denoiser for removing white Gaussian noise and the cross-entropy-optimal classifier for predicting the noise level. Our method is the first DRE-based technique that can successfully generate images beyond the MNIST dataset. Furthermore, it can output the likelihood of any input in a single forward pass, achieving state-of-the-art negative log likelihood (NLL) among methods with this property. Code is available on the project’s webpage.
@@ 1 Introduction
A classical family of methods for learning data distributions relies on the concept of density-ratio estimation (DRE) [46]. DRE techniques transform the unsupervised task of learning the distribution of data into the supervised task of learning to classify between data samples and samples from some reference distribution [15, 4, 35, 7]. These methods have attracted significant research efforts over the years [27, 14, 35, 47], particularly for their inherent capability to directly output the likelihood for any given input. However, to date, they have not succeeded in capturing the distribution of complex high-dimensional data, like natural images. Instead, their generative performance was demonstrated only on low-dimensional toy examples and on the simple MNIST handwritten digits dataset [28, 35, 7]. As illustrated in Fig. 1, while the state-of-the-art DRE method, telescoping density-ratio estimation (TRE) [35], succeeds in capturing the distribution of the MNIST dataset [28], it fails on the slightly more complex CIFAR-10 dataset [26].
@@
Our experiments shed light on the reasons why DRE methods have failed on complex high-dimensional data to date, and why CDM inherently avoids these challenges. Furthermore, we show that CDM can serve as a more accurate denoiser, in terms of MSE, than a DDM with a similar architecture. This typically translates into better FID scores. Representative generated samples are shown in Fig. 1. We hope that our approach will spark new interest in DRE methods and ultimately unlock their full potential.
Figure 1: Samples from CDMs (left) trained on CelebA $\mathbf{64\times 64}$ and on CIFAR-10, compared to samples from TRE models [35] (right) trained on MNIST and CIFAR-10. To date, DRE methods have failed to capture the distributions of complex, high-dimensional data, and have been demonstrated only on toy examples or on the simple MNIST dataset. The right pane shows results from TRE, the state-of-the-art DRE method, which fails to capture the distribution of CIFAR-10. CDM is the first DRE-based method that can successfully learn the distribution of images.
@@ 2.1 Density Ratio Estimation
Unfortunately, this method fails in practice when $p_{d}({\bm{x}})$ and $p_{n}({\bm{x}})$ differ significantly from one another, as is the case when $p_{d}({\bm{x}})$ is the distribution of images and $p_{n}({\bm{x}})$ corresponds to white Gaussian noise. This is because when training a classifier to discriminate between images and noise, the classifier can achieve very high accuracy even without learning meaningful information about images. When this point is reached, the weights of the classifier practically stop updating. Rhodes et al. [35] referred to this issue as the density-chasm problem, and suggested to overcome it by making the classification problem more difficult. To do so, their TRE method uses a sequence of distributions $p_{{\mathbf{x}}_{0}}({\bm{x}}),p_{{\mathbf{x}}_{1}}({\bm{x}}),\ldots,p_{{\mathbf{x}}_{m}}({\bm{x}})$, which are closer to one another, such that $p_{{\mathbf{x}}_{m}}({\bm{x}})$ is the reference distribution and $p_{{\mathbf{x}}_{0}}({\bm{x}})$ is the target distribution. The intermediate distributions $\{p_{{\mathbf{x}}_{i}}({\bm{x}})\}_{i=1}^{m-1}$ do not have to be known; the only requirement is that it would be possible to sample from them. For example, ${\mathbf{x}}_{i}$ can be defined as $\smash{{\mathbf{x}}_{i}=\sqrt{\bar{\alpha}_{i}}{\mathbf{x}}_{0}+\sqrt{1-\bar{\alpha}_{i}}{\mathbf{x}}_{m}}$, where ${\mathbf{x}}_{0}\sim p_{{\mathbf{x}}_{0}}$, ${\mathbf{x}}_{m}\sim p_{{\mathbf{x}}_{m}}$, and $\bar{\alpha}_{i}$ is a sequence that decreases from 1 to 0. Then, using (1), each ratio $p_{{\mathbf{x}}_{i}}({\bm{x}})/p_{{\mathbf{x}}_{i+1}}({\bm{x}})$ can be extracted by training a binary classifier to distinguish between samples from $p_{{\mathbf{x}}_{i}}({\bm{x}})$ and $p_{{\mathbf{x}}_{i+1}}({\bm{x}})$, and the ratio between the target and the reference distributions can be calculated as
@@
While this method overcomes the density-chasm-problem for each pair of consecutive distributions, it still fails in learning the distribution of datasets that are more complicated than MNIST, as illustrated in Fig. 1. This is because each ratio $p_{{\mathbf{x}}_{i}}({\bm{x}})/p_{{\mathbf{x}}_{i+1}}({\bm{x}})$ is extracted from a binary classifier trained only on inputs ${\bm{x}}$ from the distributions $p_{{\mathbf{x}}_{i}}$ and $p_{{\mathbf{x}}_{i+1}}$. For instance, the classifier producing the ratio $p_{{\mathbf{x}}_{0}}({\bm{x}})/p_{{\mathbf{x}}_{1}}({\bm{x}})$ is trained on inputs close to the real data, while the one producing $p_{{\mathbf{x}}_{m-1}}({\bm{x}})/p_{{\mathbf{x}}_{m}}({\bm{x}})$ is trained on inputs close to the reference distribution. This can lead to a mismatch between training and test time, since at inference, all the ratios are evaluated at the same input ${\bm{x}}$. Moreover, even if each individual ratio is nearly accurate, the accumulation of small errors can result in a significant overall error. Our method is also based on a classification objective, however it avoids these problems by employing an additional loss, which is based on our main result (Theorem 3.1).
@@ 2.2 Denoising Diffusion Models
DDMs [38, 19], are a class of generative models that sample from a learned target distribution by gradually denoising white Gaussian noise. More formally, DDMs generate samples by attempting to reverse a forward diffusion process with $T$ steps that starts from a data point ${\mathbf{x}}_{0}$ and evolves as ${\mathbf{x}}_{t}=\sqrt{\alpha_{t}}{\mathbf{x}}_{t-1}+\sqrt{1-\alpha_{t}}\tilde{\varepsilon}_{t}$, $t=1,\ldots,T$, where $\{\tilde{\varepsilon}_{t}\}$ are iid standard Gaussian vectors. Samples along this forward diffusion process can be equivalently expressed as
@@
and covariance $\sigma_{t}\mathbf{I}$, where $\varepsilon_{\theta}(\cdot,\cdot)$ is a neural network and $\{\sigma_{t}\}$ are fixed hyperparameters. Training is done by minimizing the ELBO loss, which reduces to a series of MSE terms,
$\mathcal{L}(\theta)=\sum_{t=1}^{T}\mathbb{E}_{{\mathbf{x}}_{0},{\mathbf{\varepsilon}}_{t}}\left[\lVert\varepsilon_{\theta}({\mathbf{x}}_{t},t)-{\mathbf{\varepsilon}}_{t}\rVert_{2}^{2}\right].$ (5)
@@ 3 Method
We are interested in a classifier for predicting t from $\tilde{{\mathbf{x}}}$. It is well known that given any sample ${\bm{x}}$ drawn from (7), the optimal such classifier (in terms of the cross-entropy loss) outputs the probability vector $(p_{{\textnormal{t}}|\tilde{{\mathbf{x}}}}(0|{\bm{x}}),p_{{\textnormal{t}}|\tilde{{\mathbf{x}}}}(1|{\bm{x}}),\ldots,p_{{\textnormal{t}}|\tilde{{\mathbf{x}}}}(T+1|{\bm{x}}))$, where $p_{{\textnormal{t}}|\tilde{{\mathbf{x}}}}(t|{\bm{x}})=\mathbb{P}({\textnormal{t}}=t|\tilde{{\mathbf{x}}}={\bm{x}})$. As we now show, the denoiser in (6) corresponds to the gradient of this classifier.
@@
Theorem 3.1 suggests that we may train a classifier and use its gradient as a denoiser according to relation (8). This paradigm is illustrated in Fig. 2. Once we have constructed a denoiser, we can apply any desired sampling method (e.g. DDPM, DDIM, etc.) to generate images from the learned distribution. However, as we show in Sec. 4.1, naively training such a classifier with the standard cross-entropy (CE) loss leads to poor results. This is because a classifier may reach a low CE loss even without learning the correct probability $p_{{\textnormal{t}}|\tilde{{\mathbf{x}}}}(t|{\bm{x}})$ for any $t$. This phenomenon can be observed in Fig. 3, which illustrates the reason that existing DRE methods fail to capture the distribution of high-dimensional complex data like images. We discuss this in more detail in Sec. 4.1.
@@
To obtain the correct probability $p_{{\textnormal{t}}|\tilde{{\mathbf{x}}}}(t|{\bm{x}})$ for any $t$, we suggest training the classifier with a combination of a CE loss on its outputs, and an MSE loss on its gradient, according to relation (8). The network’s gradient can be efficiently computed using automatic-differentiation. Following Ho et al. [19], we use the same weight for all timesteps in the MSE loss. Our full training scheme is described in Algorithm 1. Here, $f_{\theta}({\bm{x}})[t]$ denotes the model’s $t$-th logit, which serves as an approximation for $\log p_{{\textnormal{t}}|\tilde{{\mathbf{x}}}}(t|{\bm{x}})$ (up to an additive constant that cancels out in the SoftMax operation), and $\smash{F_{\theta}({\bm{x}},t)=f_{\theta}({\bm{x}})[T+1]-f_{\theta}({\bm{x}})[t]}$. The added timesteps, corresponding to entries 00 and $T+1$ of the classifier, are trained only using the CE loss. This is because the prediction of the noise is trivial when ${\textnormal{t}}=T+1$ and meaningless when ${\textnormal{t}}=0$ (since there is no noise). Importantly, this behavior is automatically achieved without any modification to the algorithm. Specifically, in line 4 of the algorithm, $F_{\theta}({\bm{x}}_{T+1},T+1)=0$, and in line 5, $\varepsilon_{\theta}({\bm{x}}_{0},0)=0$, preventing the MSE loss from updating the weights for these timesteps.
Algorithm 2 shows how to generate samples with CDM using the DDPM sampler (a similar approach can be used with other samplers). Note that each step $t$ in DDPM sampling using CDM is given by
@@ 3.1 Exact Likelihood Calculation in a Single Step
where $\mathcal{N}(\cdot;0,\mathbf{I})$ is the probability density function of a standard multivariate Gaussian distribution.
Note that the first term in (10) only depends on the pre-selected probability mass function $p_{\textnormal{t}}$ (which we choose to be uniform in our experiments), and the second term can be obtained from the $t$-th and $(T+1)$-th entries of the vector at the output of the classifier (after applying SoftMax). This implies that we can calculate the likelihood of any given image ${\bm{x}}$ w.r.t. to the density $p_{{\mathbf{x}}_{t}}$ of noisy images for any noise level $t$. In particular, $p_{{\mathbf{x}}_{0}}$ is the density of clean images. Thus, by choosing $t=0$ we can calculate the likelihood of clean images. In App. D we present a simple validation of our efficient likelihood computation by applying it to toy examples with known densities, allowing us to compute the analytical likelihood and verify that our computations align with theoretical expectations.
Figure 3: Comparison between the log probability of a noise-level classifier trained using the CE loss alone and a model trained using CE and MSE. Since the SoftMax operator is invariant to an additive factor, we subtract the maximal value from the vector (i.e., $f_{\theta}({\bm{x}}_{t})\leftarrow f_{\theta}({\bm{x}}_{t})-\max(f_{\theta}({\bm{x}}_{t}))$) for visualization. We utilize the connection we developed between the optimal classifier and the MMSE denoiser to incorporate the MSE loss in DRE training, as depicted in Algorithm 1. As evident, without considering MSE, the prediction accuracy of the classifier is limited to the vicinity of the correct label, unlike the model trained using both CE and MSE, which yields accurate predictions globally. The essence of an optimal classifier lies in its capability to predict the correct probability vector for all entries, rather than solely for the correct label. As can be seen, this necessitates the incorporation of the MSE loss.